Whiterock Corporation issued a bond issue to investors on January 2, 2016. The 20-year corporate bond
has a contract rate of 5.5%. The contract terms specified the following requirements at the end of each year
until the bonds mature:
Working capital of $2,000,000
Current ratio of 2.0
Quick ratio of 1.8
Whiterock calculated working capital, current ratio and the quick ratio based upon the following determinations:
Current Liquidity Measurements
Current Assets: Working Capital = Current Assets - Current Liabilities
Cash 1,200,000 2,948,000
Marketable securities 260,000
Accounts receivable 2,750,000
Inventories 440,000 Current Ratio
Prepaid Insurance and Rent 48,000 2.00
Intangible Assets 1,200,000
Total Current Assets 5,898,000
Quick Ratio
Current Liabilities 1.83
Accounts Payable 1,600,000
Wages and Salaries Payable 350,000
Notes Payable (Short-term) 1,000,000 2,950,000
What is the Revised Liquidity Measures:
Working Capital = Current Assets - Current Liabilities
Current Ratio = Current Assets
Current Liabilities
Quick Ratio = Quick Assets
Current Liabilities
Answer #1
Current Assets:
Cash 1,200,000
Marketable securities 260,000
Accounts receivable 2,750,000
Inventories 440,000
Prepaid Insurance and Rent 48,000
Total current assets 4,698,000
Current Liabilities
Accounts Payable 1,600,000
Wages and Salaries Payable 350,000
Notes Payable (Short-term) 1,000,000
Total current liabilities 2,950,000
Working Capital = Current Assets - Current Liabilities
4698000-2950000= 1748000
Current Ratio =
Current assets/Current Liabilities
4698000/2950000= 1.59
Quick Ratio =
(Current assets-Inventories & Prepaid assets)/Current Liabilities
(4698000-440000-48000)/2950000= 1.43